OpenStudy (anonymous):

Inverse Variation Help Please!!! The pair of points is on the graph of an inverse variation. Find the missing value. (2.4, 3) and (5, y) A) 1.44 B) 1 C) 6.25 D) 0.69

OpenStudy (godlovesme):

sorry idk this stufff :( @Nnesha help :) he/she is genius

OpenStudy (ilovecake):

Let's solve the first point and by doing so we have to find the constant of variation which y = k/x and 3 = k/2.4. Using this same constant of variation with the second point, 2.4 *3=7.2 y = k/x y = 7.2/5 y = ?

OpenStudy (anonymous):

1.44? @Ilovecake

OpenStudy (ilovecake):

yup

OpenStudy (godlovesme):

@Ilovecake u r awesome i'm sooo fanning u :)

OpenStudy (anonymous):

Thank you SO much! Can you help me with a few more?? @Ilovecake

OpenStudy (ilovecake):

yea sure

OpenStudy (ilovecake):

ok

OpenStudy (anonymous):

Find the constant of variation k for the inverse variation. Then write an equation for the inverse variation. y = 5 when x = 7.5 A) k = 15; xy = 1.5 B) k = 37.5; xy = 37.5 C) k = 37.5; 37.5y = x D) k = 2.5; y = 2.5/x

OpenStudy (anonymous):

I know the answer isn't A or D...

OpenStudy (ilovecake):

\[y = 37 \over x \] and now we can check. 5= 37.5

OpenStudy (ilovecake):

37.5 over 7.5

OpenStudy (anonymous):

I don't see that option though...

OpenStudy (anonymous):

Would it be B?

OpenStudy (ilovecake):

It would be option C.....

OpenStudy (anonymous):

C was incorrect... but can you help me with some more?

OpenStudy (ilovecake):

it was incorrect???

OpenStudy (ilovecake):

sure I will help you more